Unit Conversions:
Convert 7.71 m/s^2 to km/h^2

(Correct Answer: 1 km/h^2)

Please show all of your work!!! Thank you
Answer #1

Given : acceleration: 7.71 m/s2

Goal : Unit conversion to km/h2

Solution:

We know that , 1 km = 1000 m

And 1 m = (1/1000) km

So , converting m/s2 to km/s2 , we get :

7.71 m/s2 = (7.71/1000) km/s2 = 0.00771 km/s2

Now , we know that , 1 h = 60 min = 3600 s

1s = (1/3600) h

Now converting , km/s2 to km/h2 , we get:

0.00771 km/s2 = [0.00771÷(1/3600)] km/h2

= 99921.6 km/h2

Answer :    7.71 m/s2 = 99921.6 km/h2

Note :   1 km/h2 = 7.71 × 10-5 m/s2

And 1 m/s2 = 12960 km/h2